19:30: telescope open, flats taken, we notice a strange pattern on LUCI1 detector, probably it is due to consensation on LUCI entrance window.

NGC2419

20:20: preset sent as mag 14. Problems with timeout when preset is sent through LUCI. We are also facing into the wind, shells are failing.

LAT_20160313.png

Hip49046

20:50: still having problems with presets, timeout occurring...

lgsw.tools.patrol_lgs_jitter: procedure to acquire fast frames from patrol cameras.

LBFC: it seems that no astigmatism is present, corrector at 76 steps, focus at -80k steps:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160313_044236 0, 0, 0 NA NA 1.43133 F F LBFC opt
20160313_044254 0, 0, 0 NA NA 1.56875 F F LBFC opt
20160313_044310 0, 0, 0 NA NA 1.61722 F F LBFC opt
20160313_044328 0, 0, 0 NA NA 1.68334 F F LBFC opt
20160313_044345 0, 0, 0 NA NA 1.74727 F F LBFC opt
20160313_044402 0, 0, 0 NA NA 1.8105 F F LBFC opt
20160313_lbfc.png

Then we acquire the IM between jitter and LAS pupil mirror.

New pupil positions on-sky:

Pupil X Y
Blue 340 420
Yellow 360 440
Red 525 410
PBL also with BCU: it fades when closing loop (we're using PYR as TT sensor).

23:30: PBL wiht PCD: it is shutdown and it does not restart!

00:00: AOS timeout fixed (see Josè and Steve's e-mails of 13/03/2016: "Changes made to aos.conf")

NGC4449

00:20: preset as mag 14.6, mag 15.7 on PYR (FW1= 600-1000)

SX and DX PCD are swapped

NGC5921

3:20: preset as mag 16.5, mag 17 on PYR (FW1= 600-1000)

We could close the AO loop for at least 1min...

LBFC: corrector @ 530 steps, focus @ -150k.

20160313_095200_lbc.png

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160313_100858 1.5,0.4,0.35 184.168 20151212_233600 1.24343 T F
20160313_100924 1.5,0.4,0.35 193.87 20151212_233600 1.10699 T F
5:20: BCU voltage trimmed to 3.485V

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160313_113257 1.5,0.25,0.1 255.214 20151212_233600 0.867335 T F
20160313_113320 1.5,0.25,0.1 251.279 20151212_233600 0.90884 T F
20160313_114243 0.2,0.1, 0 1300.81 20151212_233600 0.976432 T F
20160313_114302 0.2,0.1, 0 1326.88 20151212_233600 0.909394 T F
20160313_114319 0.2,0.1, 0 488.533 20151212_233600 0.965491 T F
20160313_114402 0.2,0.1, 0 457.264 20151212_233600 1.07509 T F
20160313_114418 0.2,0.1, 0 514.632 20151212_233600 1.09358 T F
20160313_114434 0.2,0.1, 0 444.984 20151212_233600 1.085 T F
20160313_114511 0.2,0.1, 0 607.089 20151212_233600 1.04035 T F
20160313_114526 0.2,0.1, 0 528.837 20151212_233600 0.996058 T F
20160313_114545 0.2,0.1, 0 396.179 20151212_233600 1.04953 T F
20160313_114617 0.2,0.1, 0 486.944 20151212_233600 0.971762 T F
20160313_120034 1.5,0.4,0.3 265.277 20151212_233600 -1 T F
jitterResidualWithAndWithoutVibComp.png

LBFC:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160313_123034 1.5, 0, 0 348.983 20151212_233600 -1 F F LBFC opt
20160313_123053 1.5, 0, 0 352.997 20151212_233600 -1 F F LBFC opt
20160313_123110 1.5, 0, 0 345.053 20151212_233600 -1 F F LBFC opt
20160313_123129 1.5, 0, 0 341.813 20151212_233600 -1 F F LBFC opt
20160313_123148 1.5, 0, 0 344.79 20151212_233600 -1 F F LBFC opt
20160313_123206 1.5, 0, 0 342.503 20151212_233600 -1 F F LBFC opt
-- MarcoBonaglia - 22 Mar 2016
Topic revision: r1 - 22 Mar 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback